OpenStudy (anonymous):

1/5(P-5)= 3/5P + 1/10P + 1 I am studying for a test in college intermediate algebra. You may be seeing me post questions because I am very confused on some of these question types. Please do not just give me the answer, explain the best you can to help me understand how to work it out for my test.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Wait, how did you get 3/10P+1? Sorry, like i said its a little confusing for me.

OpenStudy (anonymous):

i bet the \(p-5\) is not in the denominator

OpenStudy (anonymous):

No. it's the same question here i'll draw it in.

OpenStudy (anonymous):

\[\frac{1}{5}(p-5)=\frac{3}{5}p+\frac{1}{10}p\]

OpenStudy (anonymous):

Thank you satellite not sure how to draw it in.

OpenStudy (anonymous):

\[\frac{1}{5}(p-5)=\frac{3}{5}p+\frac{1}{10}p+1\]

OpenStudy (anonymous):

multiply both sides by 10 \[2(p-5)=6p+p+10\] and go from there

Nnesha (nnesha):

oky nvm sorry about that i typed it wrong

OpenStudy (anonymous):

I got p= -2 however, my book says -4?

OpenStudy (anonymous):

?

Nnesha (nnesha):

yes -2 is not right

Nnesha (nnesha):

can you show ur work so i can figure out where u did a mistake :)